# Structuring Multiple Simple Cycle Reservoirs with Particle Swarm Optimization

### v1

@@ II Notion of MSCR
In this section, we introduce the notion of reservoir systems, Simple Cycle Reservoirs (SCR) and Multi-SCR (MSCR), focused throughout the paper.
@@
The contractiveness of W and the uniform boundedness of input stream imply that the images $\mbox{\bf{x}}\in\mathbb{R}^{n}$ of the inputs $\mbox{\bf{c}}\in(\mathbb{R}^{m})^{\mathbb{Z}_{-}}$ under the linear reservoir system live in a compact space $X\subset{\mathbb{R}}^{n}$. With slight abuse of mathematical terminology we call $X$ a state space.
@@
A reservoir system $R=\left(\mbox{\bf{W}},\mbox{\bf{W}}_{in},f,h\right)$ with dimensions $(n,m,d)$ is called a Simple Cycle Reservoir (SCR)  [^2] We note that the assumption on the aperiodicity of the sign pattern in $V$ is not required for this study if
@@
The reachability matrix $\mathcal{R}$ can be computed from the connectivity matrix A using Floyd–Warshall algorithm [28] with $O(k^{3})$.
@@
For the rest of the paper, we will consider reservoir systems operating on univariate input, i.e. the input dimension is $m=1$. The input and target are denoted by $\{u_{t}\}_{t\in\mathbb{Z}_{-}}\subset\mathbb{R}$ and $\{\mbox{\bf{y}}(t)\}_{t\in\mathbb{Z}_{-}}\subset\mathbb{R}$, respectively. The spectral radius of the $i^{th}$ vertex-encoder $R^{(i)}$ in a Multi-Simple Cycle Reservoir of order $k$ will be denoted by $\rho^{(i)}$. In this work, we consider the data flow in the MSCR follows rules listed as follows:
@@ III Particle Swarm Optimization
Particle Swarm Optimization (PSO) [29] on Euclidean space $\mathbb{R}^{n}$ is a population-based meta-heuristic optimization algorithm. PSO is a zeroth order algorithm motivated by swarm intelligence. PSO makes minimal assumptions on the search space – it does not require the objective function to be differentiable and only pairwise addition and negation is required for the search space.
@@
  • Inertial Coefficient ($w^{\ell}$): weight of the inertia component, a predefined real-valued function on the iteration counter $\ell$
@@ IV-A Model selection
We denote an MSCR optimized via PSO as MSCR-PSO. For the optimization of D and A, the particle positions at the $l$-th iteration are converted to binary values using the Heaviside step function.
@@
  • MSCR-GA builds upon recent work using a GA-based method to optimize only the network topology of an MRESN with pre-determined input and inter-scaling factors. In this study, we replaced the randomly initialized vertex-encoders with SCR vertex-encoder to construct MSCR-GA. This comparison aims to measure the performance gains achieved by simultaneously optimizing scaling and topology factors in MSCR-PSO.
@@
For the numerical experiments, we introduce a bias term in the dynamics of MSCR such as in [23]. In particular, for each vertex-encoder of an MSCR, denoted by $R^{(i)}:=(\mbox{\bf{W}}^{(i)},{\mbox{\bf{s}}}_{i}\cdot\mathcal{W}_{in}^{(i)},f^{(i)},h^{(i)})$, a small bias vector ${\mbox{\bf{b}}}^{(i)}$ is introduced to the state evolution. In particular, the dynamics of reservoir system now reads:
@@ IV-B Parameter settings
We fixed the dimension of state space for an SCR at 1000. For an MSCR model, we set the maximal order $k^{max}$ to 10, and the mapping space of each vertex-encoder to $n=100$. We set the spectral radius for each vertex-encoder to $\rho=0.95$ and used ridge regression with a regularization factor 1E-4 for training readout function $h(\cdot)$. Note that the hyperparameters we searched in the experiments are initialization methods of $\mathbf{V}_{in}^{i}$ and $\mathbf{V}^{j}_{i}$ for $i=1,...,k$, and types of activation function $f$. Coupling matrices $\mathbf{V}_{in}^{i}$ and $\mathbf{V}^{j}_{i}$ are initialized by either the binary expansion of $\pi$ of the appropriate size or sampled randomly using Bernoulli trails over $\left\{-1,1\right\}$. The activation function was selected from the identity function and the hyperbolic tangent function. We recorded the average prediction performances across ten trials when the Bernoulli-based initialization method was adopted.
@@ IV-C Datasets and task settings
For all datasets, we partitioned the total data length into the lengths of the washout, training, validation, and test sets, which were set at $100$, $1000$, $1000$, and $1000$ respectively. The glimpses of datasets are shown in Fig. 2.
@@ IV-E Results
Tables II-IV present the models that achieved the best performance on the three datasets, along with their corresponding optimal hyperparameter settings, which includes the choice of distribution and activation function.
Overall, MSCR-PSO out-perform a single SCR across all three datasets, which indicates that an MSCR with several vertex-encoders organized by an optimized network topology is more effective than a single large SCR.
@@
Furthermore, MSCR-PSO demonstrates substantial improvements of approximately 85.21%, 58.66%, and 1.67% compared to MSCR-GA on MG-17, NARMA-10, and MSTSN respectively. These results indicate that the PSO-based network topology optimization method is more effective than the GA-based network topology for an MSCR system on time-series prediction tasks.
@@ V Discussion
Across the three benchmark datasets, we found that MSCR-PSO achieves competitive predictive performance while utilizing a lower-dimensional state space compared to both SCR and MSCR with pre-determined network topologies. The significance of this work lies in the advancements of MSCR-PSO over hand-crafted MSCR, which pave the way for systemic optimization of task-specific multi-reservoir systems.
Furthermore, since PSO operates on continuous values, experimental results empirically demonstrate its superiority over existing GA-based optimization methods for optimizing multi-reservoir systems.
@@ Acknowledgment
This work was partly supported by JSPS KAKENHI Grant Number JP23K28154 (GT), JST CREST Grant Numbers JPMJCR19K2 (ZL, KF, GT), JPMJCR24R2 (GT), and JST Moonshot R&D Grant Number JPMJMS2021 (ZL, KF, KA, GT), Cross-ministerial Strategic Innovation Promotion Program (SIP), the 3rd period of SIP, Grant Numbers JPJ012207 (KA), JPJ012425 (KA), Institute of AI and Beyond of UTokyo.

### v2

@@ II Notion of MSCR
In this section, we introduce the notion of reservoir systems, Simple Cycle Reservoirs (SCR), and Multi-SCR (MSCR), which are focused on throughout the paper.
@@
The contractiveness of W and the uniform boundedness of the input stream imply that the images $\mbox{\bf{x}}\in\mathbb{R}^{n}$ of the inputs $\mbox{\bf{c}}\in(\mathbb{R}^{m})^{\mathbb{Z}_{-}}$ under the linear reservoir system live in a compact space $X\subset{\mathbb{R}}^{n}$. With slight abuse of mathematical terminology, we call $X$ a state space.
@@
A reservoir system $R=\left(\mbox{\bf{W}},\mbox{\bf{W}}_{in},f,h\right)$ with dimensions $(n,m,d)$ is called a Simple Cycle Reservoir (SCR)  [^2] We note that the assumption on the aperiodicity of the sign pattern in $V$ is not required for this study. if
@@
The reachability matrix $\mathcal{R}$ can be computed from the connectivity matrix A using the Floyd-Warshall algorithm [28] with $O(k^{3})$.
@@
For the rest of the paper, we will consider reservoir systems operating on univariate input, i.e., the input dimension is $m=1$. The input and target are denoted by $\{u_{t}\}_{t\in\mathbb{Z}_{-}}\subset\mathbb{R}$ and $\{\mbox{\bf{y}}(t)\}_{t\in\mathbb{Z}_{-}}\subset\mathbb{R}$, respectively. The spectral radius of the $i^{th}$ vertex-encoder $R^{(i)}$ in a Multi-Simple Cycle Reservoir of order $k$ will be denoted by $\rho^{(i)}$. In this work, we consider the data flow in the MSCR follows the rules listed as follows:
@@ III Particle Swarm Optimization
Particle Swarm Optimization (PSO) [29] on Euclidean space $\mathbb{R}^{n}$ is a population-based meta-heuristic optimization algorithm. PSO is a zeroth-order algorithm motivated by swarm intelligence. PSO makes minimal assumptions on the search space – it does not require the objective function to be differentiable, and only pairwise addition and negation are required for the search space.
@@
  • Inertial Coefficient ($w^{\ell}$): weight of the inertia component, a predefined real-valued function on the iteration counter $\ell$.
@@ IV-A Model selection
We denote an MSCR optimized via PSO as MSCR-PSO. For the optimization of d and A, the particle positions at the $l$-th iteration is converted to binary values by mapping all positive values as one and negative values as zero.
@@
  • MSCR-GA builds upon recent work using a GA-based method to optimize only the network topology of an MRESN with pre-determined input and inter-scaling factors. In this study, we replaced the randomly initialized vertex-encoders with the SCR vertex-encoder to construct MSCR-GA. This comparison aims to measure the performance gains achieved by simultaneously optimizing scaling and topology factors in MSCR-PSO.
@@
For the numerical experiments, we introduce a bias term in the dynamics of MSCR, such as in [23]. In particular, for each vertex-encoder of an MSCR, denoted by $R^{(i)}:=(\mbox{\bf{W}}^{(i)},{\mbox{\bf{s}}}_{i}\cdot\mathcal{W}_{in}^{(i)},f^{(i)},h^{(i)})$, a small bias vector ${\mbox{\bf{b}}}^{(i)}$ is introduced to the state evolution. In particular, the dynamics of the reservoir system now reads:
@@ IV-B Parameter settings
We fixed the dimension of the state space for an SCR at 1000. For an MSCR model, we set the maximal order $k^{max}$ to 10, and the mapping space of each vertex-encoder to $n=100$. We set the spectral radius for each vertex-encoder to $\rho=0.95$ and used ridge regression with a regularization factor 1E-4 for training the readout function $h(\cdot)$. Note that the hyperparameters we searched in the experiments are initialization methods of $\mathbf{V}_{in}^{i}$ and $\mathbf{V}^{j}_{i}$ for $i=1,...,k$, and types of activation function $f$. Coupling matrices $\mathbf{V}_{in}^{i}$ and $\mathbf{V}^{j}_{i}$ are initialized by either the binary expansion of $\pi$ of the appropriate size or sampled randomly using Bernoulli trails over $\left\{-1,1\right\}$. The activation function was selected from the identity function and the hyperbolic tangent function. We recorded the average prediction performances across ten trials when the Bernoulli-based initialization method was adopted.
@@ IV-C Datasets and task settings
For all datasets, we partitioned the total data length into the lengths of the washout, training, validation, and test sets, which were set at $100$, $1000$, $1000$, and $1000$, respectively. The glimpses of datasets are shown in Fig. 2.
@@ IV-E Results
Tables II-IV present the models that achieved the best performance on the three datasets, along with their corresponding optimal hyperparameter settings, which include the choice of distribution and activation function.
Overall, MSCR-PSO outperforms a single SCR across all three datasets, which indicates that an MSCR with several vertex-encoders organized by an optimized network topology is more effective than a single large SCR.
@@
Furthermore, MSCR-PSO demonstrates substantial improvements of approximately 85.21%, 58.66%, and 1.67% compared to MSCR-GA on MG-17, NARMA-10, and MSTSN, respectively. These results indicate that the PSO-based network topology optimization method is more effective than the GA-based network topology for an MSCR system on time-series prediction tasks.
@@ V Discussion
Across the three benchmark datasets, we found that MSCR-PSO achieves competitive predictive performance while utilizing a lower-dimensional state space compared to both SCR and MSCR with pre-determined network topologies. The significance of this work lies in the advancements of MSCR-PSO over hand-crafted MSCR, which paves the way for systemic optimization of task-specific multi-reservoir systems.
Furthermore, since PSO operates on continuous values, it can achieve both of the following simultaneously:
  • direct optimization of tune real‑valued entries, such as H and s, and
  • optimize the binary entries of both A and d via specific transformation functions.
In contrast, GA relies on discrete encoding and is inherently limited to categorical or binary search. This flexibility of operating on continuous values makes PSO particularly well‑suited for constructing MSCR. Experimental results also empirically demonstrate its superiority over existing GA-based optimization methods for optimizing multi-reservoir systems.
@@ Acknowledgment
This work was partly supported by JSPS KAKENHI Grant Number JP23K28154 (GT), JST CREST Grant Numbers JPMJCR19K2 (ZL, KF, GT), JPMJCR24R2 (GT), and JST Moonshot R&D Grant Number JPMJMS2021 (ZL, KF, KA, GT), Cross-ministerial Strategic Innovation Promotion Program (SIP), the 3rd period of SIP, Grant Numbers JPJ012207 (KA), JPJ012425 (KA, KF), Institute of AI and Beyond of UTokyo (KA), the World Premier International Research Center Initiative (WPI), MEXT, Japan.